GhostlyDawnM:

Which BEST describes the system of equations? x = y − 3 −4x = 12 − 4y A) Consistent B) Inconsistent C) Consistent and Dependent D) Consistent and Independent

GhostlyDawnM:

it is B) inconsistent

ramen:

what do you think so far?

ramen:

try finding the definitions first

ramen:

then you can check your answer

Mercury:

old question but will respond so this can be closed I disagree with inconsistent for the reasons I will discuss below: x = y − 3 −4x = 12 − 4y notice how, if you multiply the first equation by -4, and re-arrange accordingly, both equations are equal. one of the other answer choices better describes a system with two equivalent equations (two overlapping graphs)